# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 5.1 Preliminaries
(5.3) $\begin{array}[]{lllll}\displaystyle\widetilde{a}_{\boldsymbol{\sigma}_{h},p_{h}}({\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}},{\boldsymbol{\psi}}_{h})+\widetilde{b}({\boldsymbol{\psi}}_{h},\chi_{p_{h}})+\widetilde{b}({\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}},q_{h})-\widetilde{c}(\chi_{p},q_{h})+\widetilde{a}_{\boldsymbol{\sigma}_{h},p_{h}}(\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}},{\boldsymbol{\psi}}_{h})\\ \displaystyle\qquad=-\widetilde{b}({\boldsymbol{\psi}}_{h},\xi_{p})-\widetilde{b}(\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}},q_{h})+\widetilde{c}(\xi_{p},q_{h})-\int_{\Omega}\Big{(}\kappa(\boldsymbol{\sigma},p)^{-1}-\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}\Big{)}\,{\boldsymbol{\varphi}}\cdot{\boldsymbol{\psi}}_{h}+c(\texttt{e}_{\boldsymbol{\sigma}},q_{h}).\end{array}$
@@ 5.2 Derivation of Céa estimates
(5.4) $\|{\boldsymbol{\chi}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{u}}\|_{0,\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{\rho}}\|_{0,\Omega}\leq C_{3}^{*}(\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{u}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{p}\|_{0,\Omega})+C_{4}^{*}\gamma\|\chi_{p}\|_{0,\Omega}.$
Proof 5.2.
@@
$\|{\boldsymbol{\chi}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{u}}\|_{0,\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{\rho}}\|_{0,\Omega}\leq(C_{1}^{*}+C_{2}^{*})\sup_{\boldsymbol{0}\neq(\boldsymbol{\tau}_{h},\boldsymbol{v}_{h},{\boldsymbol{\psi}}_{h})\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h}}\!\!\!\frac{a({\boldsymbol{\chi}}_{\boldsymbol{\sigma}},\boldsymbol{\tau}_{h})+b(\boldsymbol{\tau}_{h},({\boldsymbol{\chi}}_{\boldsymbol{u}},{\boldsymbol{\chi}}_{\boldsymbol{\rho}}))+b({\boldsymbol{\chi}}_{\boldsymbol{\sigma}},(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h}))}{\|\boldsymbol{\tau}_{h}\|_{\mathbf{div};\Omega}+\|\boldsymbol{v}_{h}\|_{0,\Omega}+\|{\boldsymbol{\eta}}_{h}\|_{0,\Omega}},$
with $C_{1}^{*},\,C_{2}^{*}>0$ independent of $h$, are the discrete version of the constants $C_{1},\,C_{2}$ defined in (3.8). Then, combining the last inequality with (5.2), and the continuity properties of $a$ and $b$ (see (2.4a)), we obtain
$\begin{array}[]{ll}\|{\boldsymbol{\chi}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{u}}\|_{0,\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{\rho}}\|_{0,\Omega}\\ \displaystyle\qquad\leq(C_{1}^{*}+C_{2}^{*})(\dfrac{1}{\mu}\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{u}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\gamma\|\chi_{p}\|_{0,\Omega}+\gamma\|\xi_{p}\|_{0,\Omega}),\end{array}$
which implies (5.4) with $C_{3}^{*}:=(C_{1}^{*}+C_{2}^{*})(\frac{1}{\mu}+1+\gamma)$ and $C_{4}^{*}:=C_{1}^{*}+C_{2}^{*}$ , and concludes the proof.
@@
(5.5) $\begin{array}[]{ll}\displaystyle\|{\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|{\boldsymbol{\chi}}_{p}\|_{0,\Omega}\leq\widetilde{C}^{*}_{5}\,(\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega})\\[4.30554pt] \qquad+\widetilde{C}^{*}_{6}\big{(}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega}+\gamma\,r\,)\|\chi_{p}\|_{0,\Omega}+\gamma\|{\boldsymbol{\chi}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}\big{)}.\end{array}$
Proof 5.4.
@@
$\|{\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}\leq 2\widetilde{C}^{*}\sup_{\boldsymbol{0}\neq({\boldsymbol{\psi}}_{h},q_{h})\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}}\frac{\widetilde{a}_{\boldsymbol{\sigma}_{h},p_{h}}({\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}},{\boldsymbol{\psi}}_{h})+b({\boldsymbol{\psi}}_{h},\chi_{p})+b({\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}},q_{h})-\widetilde{c}(\chi_{p},q_{h})}{\|{\boldsymbol{\psi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}+\|q_{h}\|_{0,\Omega}},$
@@
with $\widetilde{C}^{*}$ defined as in (4.11). Then, using the equation (5.3), the bound properties of $\widetilde{a}$, $\widetilde{b}$ and $\widetilde{c}$ (see (3.2a) and (2.4c)), and the second assumption for $\kappa$ (cf. (2.3)), we obtain
$\|{\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}$
@@
$\leq 2\widetilde{C}^{*}(C_{\widetilde{a}}\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\widetilde{\gamma}\|\xi_{p}\|_{0,\Omega}+\|\kappa(\boldsymbol{\sigma},p)^{-1}-\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}\|_{\mathbb{L}^{\infty}(\Omega)}\|{\boldsymbol{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\gamma\|\texttt{e}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega})$
@@
$\leq 2\widetilde{C}^{*}(C_{\widetilde{a}}\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\widetilde{\gamma}\|\xi_{p}\|_{0,\Omega}+\kappa_{2}\|\texttt{e}_{p}\|_{0,\Omega}\|{\boldsymbol{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\gamma\|\texttt{e}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}),$
@@
hence, using the fact that ${\boldsymbol{\varphi}}$ satisfies (3.16) and the error decomposition (5.1), we have
$\|{\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}$ $\leq 2\widetilde{C}^{*}(C_{\widetilde{a}}\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\widetilde{\gamma}\|\xi_{p}\|_{0,\Omega}+\kappa_{2}\|\xi_{p}\|_{0,\Omega}\|{\boldsymbol{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\gamma\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega})$
@@
$\qquad+2\widetilde{C}^{*}(\kappa_{2}\|\chi_{p}\|_{0,\Omega}\|{\boldsymbol{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\gamma\|{\boldsymbol{\chi}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega})$
$\leq\widetilde{C}^{*}_{5}\,(\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega})$
$\qquad+2\widetilde{C}^{*}\big{(}\kappa_{2}C(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega}+\gamma\,r\,)\|\chi_{p}\|_{0,\Omega}+\gamma\|{\boldsymbol{\chi}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}\big{)},$
@@
the last equation implies (5.5), with $\widetilde{C}^{*}_{5}:=2\widetilde{C}^{*}\big{(}C_{\widetilde{a}}+1+\widetilde{\gamma}+\gamma+\kappa_{2}C(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega}+\gamma\,r\,)\big{)}$ and $\widetilde{C}^{*}_{6}:=2\widetilde{C}^{*}(\kappa_{2}C+1)$, and concludes the proof.
@@
(5.6) $(C_{4}^{*}+\widetilde{C}^{*}_{6}+\widetilde{C}^{*}_{6}r)\gamma+\widetilde{C}^{*}_{6}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega})\leq\dfrac{1}{2},$
with $C_{4}^{*}$ and $\widetilde{C}^{*}_{6}$ being the constants in Lemmas 5.1 and 5.3. Furthermore, assume that the hypotheses of Theorem 3.12 and Theorem 4.1 hold. Let $(\boldsymbol{\sigma},\boldsymbol{u},\boldsymbol{\rho},{\boldsymbol{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$ and $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\boldsymbol{\rho}_{h},{\boldsymbol{\varphi}}_{h},p_{h})\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h}\times\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$ be the unique solutions of (2.2) and (4.6), respectively. Then, there exists $C_{\mathrm{C\acute{e}a}}>0$, independent of $h$, such that
(5.7) $\begin{array}[]{ll}\displaystyle\|\texttt{e}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{\boldsymbol{u}}\|_{0,\Omega}+\|\texttt{e}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}\\[4.30554pt] \displaystyle\qquad\leq C_{\mathrm{C\acute{e}a}}\,\mathrm{dist}\,\big{(}(\boldsymbol{\sigma},\boldsymbol{u},\boldsymbol{\rho},{\boldsymbol{\varphi}},p),\,\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h}\times\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}\big{)}.\end{array}$
Proof 5.6.
@@
$\begin{array}[]{ll}\displaystyle\|{\boldsymbol{\chi}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{u}}\|_{0,\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|{\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}\leq C_{3}^{*}(\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{u}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\xi_{p}\|_{0,\Omega})\\[4.30554pt] \qquad\displaystyle+\widetilde{C}^{*}_{5}\,(\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega})+\dfrac{1}{2}\|\chi_{p}\|_{0,\Omega}+\dfrac{1}{2}\|{\boldsymbol{\chi}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}.\end{array}$

### v2

@@ 5.1 Preliminaries
$\begin{array}[]{lllll}\displaystyle\widetilde{a}_{\bm{\sigma}_{h},p_{h}}({\bm{\chi}}_{{\bm{\varphi}}},{\bm{\psi}}_{h})+\widetilde{b}({\bm{\psi}}_{h},\chi_{p_{h}})+\widetilde{b}({\bm{\chi}}_{{\bm{\varphi}}},q_{h})-\widetilde{c}(\chi_{p},q_{h})+\widetilde{a}_{\bm{\sigma}_{h},p_{h}}(\bm{\xi}_{{\bm{\varphi}}},{\bm{\psi}}_{h})\\ \displaystyle\qquad=-\widetilde{b}({\bm{\psi}}_{h},\xi_{p})-\widetilde{b}(\bm{\xi}_{{\bm{\varphi}}},q_{h})+\widetilde{c}(\xi_{p},q_{h})-\int_{\Omega}\Big{(}{\bm{\kappa}}(\bm{\sigma},p)^{-1}-{\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}\Big{)}\,{\bm{\varphi}}\cdot{\bm{\psi}}_{h}+c(\texttt{e}_{\bm{\sigma}},q_{h}).\end{array}$ (5.3)
@@ 5.2 Derivation of Céa estimates
$\|{\bm{\chi}}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|{\bm{\chi}}_{{\bm{u}}}\|_{0,\Omega}+\|{\bm{\chi}}_{\bm{\rho}}\|_{0,\Omega}\leq C_{3}^{*}(\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\bm{\xi}_{{\bm{u}}}\|_{0,\Omega}+\|\bm{\xi}_{\bm{\rho}}\|_{0,\Omega}+\|\bm{\xi}_{p}\|_{0,\Omega})+C_{4}^{*}\gamma\|\chi_{p}\|_{0,\Omega}.$ (5.4)
Proof.
@@
$\|{\bm{\chi}}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|{\bm{\chi}}_{{\bm{u}}}\|_{0,\Omega}+\|{\bm{\chi}}_{\bm{\rho}}\|_{0,\Omega}$
$\quad\leq(C_{1}^{*}+C_{2}^{*})\sup_{\bm{0}\neq(\bm{\tau}_{h},\mathbf{v}_{h},{\bm{\psi}}_{h})\in\mathbb{H}^{\bm{\sigma}}_{h}\times\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h}}\!\!\!\frac{a({\bm{\chi}}_{\bm{\sigma}},\bm{\tau}_{h})+b(\bm{\tau}_{h},({\bm{\chi}}_{{\bm{u}}},{\bm{\chi}}_{\bm{\rho}}))+b({\bm{\chi}}_{\bm{\sigma}},(\mathbf{v}_{h},{\bm{\eta}}_{h}))}{\|\bm{\tau}_{h}\|_{\mathbf{div};\Omega}+\|\mathbf{v}_{h}\|_{0,\Omega}+\|{\bm{\eta}}_{h}\|_{0,\Omega}},$
where $C_{1}^{*},\,C_{2}^{*}>0$ independent of $h$ are the discrete version of the constants $C_{1},\,C_{2}$ defined in (3.8). Then, combining the last inequality with (5.2), and the continuity properties of $a$ and $b$ (see (2.4a)), we obtain
$\begin{array}[]{ll}\|{\bm{\chi}}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|{\bm{\chi}}_{{\bm{u}}}\|_{0,\Omega}+\|{\bm{\chi}}_{\bm{\rho}}\|_{0,\Omega}\\ \displaystyle\qquad\leq(C_{1}^{*}+C_{2}^{*})(\dfrac{1}{\mu}\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\bm{\xi}_{{\bm{u}}}\|_{0,\Omega}+\|\bm{\xi}_{\bm{\rho}}\|_{0,\Omega}+\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\gamma\|\chi_{p}\|_{0,\Omega}+\gamma\|\xi_{p}\|_{0,\Omega}),\end{array}$
which implies (5.4) with $C_{3}^{*}:=(C_{1}^{*}+C_{2}^{*})(\frac{1}{\mu}+1+\gamma)$ and $C_{4}^{*}:=C_{1}^{*}+C_{2}^{*}$. Note also, that the error estimate is robust with respect to $\lambda$. ∎
@@
$\begin{array}[]{ll}\displaystyle\|{\bm{\chi}}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|{\bm{\chi}}_{p}\|_{0,\Omega}\leq\widetilde{C}^{*}_{5}\,(\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega})\\[4.30554pt] \qquad+\widetilde{C}^{*}_{6}\big{(}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega}+\gamma\,r\,)\|\chi_{p}\|_{0,\Omega}+\gamma\|{\bm{\chi}}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}\big{)}.\end{array}$ (5.5)
Proof.
@@
$\|{\bm{\chi}}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}\leq 2\widetilde{C}^{*}\sup_{\bm{0}\neq({\bm{\psi}}_{h},q_{h})\in\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}}\frac{\widetilde{a}_{\bm{\sigma}_{h},p_{h}}({\bm{\chi}}_{{\bm{\varphi}}},{\bm{\psi}}_{h})+b({\bm{\psi}}_{h},\chi_{p})+b({\bm{\chi}}_{{\bm{\varphi}}},q_{h})-\widetilde{c}(\chi_{p},q_{h})}{\|{\bm{\psi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}+\|q_{h}\|_{0,\Omega}},$
@@
with $\widetilde{C}^{*}$ defined as in (4.9). Then, using the equation (5.3), the bound properties of $\widetilde{a}$, $\widetilde{b}$ and $\widetilde{c}$ (see (3.2a) and (2.4c)), and the second assumption for ${\bm{\kappa}}$ (cf. (2.3)), we obtain
$\|{\bm{\chi}}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}$
$\leq 2\widetilde{C}^{*}(C_{\widetilde{a}}\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\widetilde{\gamma}\|\xi_{p}\|_{0,\Omega}$
@@
$\qquad\quad+\|{\bm{\kappa}}(\bm{\sigma},p)^{-1}-{\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}\|_{\mathbb{L}^{\infty}(\Omega)}\|{\bm{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\gamma\|\texttt{e}_{\bm{\sigma}}\|_{\mathbf{div};\Omega})$
@@
$\leq 2\widetilde{C}^{*}(C_{\widetilde{a}}\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\widetilde{\gamma}\|\xi_{p}\|_{0,\Omega}+\kappa_{2}\|\texttt{e}_{p}\|_{0,\Omega}\|{\bm{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\gamma\|\texttt{e}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}),$
@@
hence, using the fact that ${\bm{\varphi}}$ satisfies (3.8) and the error decomposition (5.1), we have
$\|{\bm{\chi}}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}$ $\leq 2\widetilde{C}^{*}(C_{\widetilde{a}}\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\widetilde{\gamma}\|\xi_{p}\|_{0,\Omega}+\kappa_{2}\|\xi_{p}\|_{0,\Omega}\|{\bm{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}$
@@
$\qquad\qquad+\gamma\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega})+2\widetilde{C}^{*}(\kappa_{2}\|\chi_{p}\|_{0,\Omega}\|{\bm{\varphi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\gamma\|{\bm{\chi}}_{\bm{\sigma}}\|_{\mathbf{div};\Omega})$
$\leq\widetilde{C}^{*}_{5}\,(\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega})$
$\qquad+2\widetilde{C}^{*}\big{(}\kappa_{2}C(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega}+\gamma\,r\,)\|\chi_{p}\|_{0,\Omega}$
$\qquad\qquad+\gamma\|{\bm{\chi}}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}\big{)},$
@@
the last equation implies (5.5), with $\widetilde{C}^{*}_{5}:=2\widetilde{C}^{*}\big{(}C_{\widetilde{a}}+1+\widetilde{\gamma}+\gamma+\kappa_{2}C(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega}+\gamma\,r\,)\big{)}$ and $\widetilde{C}^{*}_{6}:=2\widetilde{C}^{*}(\kappa_{2}C+1)$, and concludes the proof. ∎
@@
$(C_{4}^{*}+\widetilde{C}^{*}_{6}+\widetilde{C}^{*}_{6}r)\gamma+\widetilde{C}^{*}_{6}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega})\leq\dfrac{1}{2},$ (5.6)
with $C_{4}^{*}$ and $\widetilde{C}^{*}_{6}$ being the constants in Lemmas 5.1 and 5.2. Furthermore, assume that the hypotheses of Theorem 3.8 and Theorem 4.1 hold. Let $(\bm{\sigma},{\bm{u}},\bm{\rho},{\bm{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$ and $(\bm{\sigma}_{h},{\bm{u}}_{h},\bm{\rho}_{h},{\bm{\varphi}}_{h},p_{h})\in\mathbb{H}^{\bm{\sigma}}_{h}\times\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h}\times\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$ be the unique solutions of (2.2) and (4.5), respectively. Then, there exists $C_{\mathrm{C\acute{e}a}}>0$, independent of $h$, such that
$\begin{array}[]{ll}\displaystyle\|\texttt{e}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{{\bm{u}}}\|_{0,\Omega}+\|\texttt{e}_{\bm{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}\\[4.30554pt] \displaystyle\qquad\leq C_{\mathrm{C\acute{e}a}}\,\mathrm{dist}\,\big{(}(\bm{\sigma},{\bm{u}},\bm{\rho},{\bm{\varphi}},p),\,\mathbb{H}^{\bm{\sigma}}_{h}\times\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h}\times\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}\big{)}.\end{array}$ (5.7)
Proof.
@@
$\|{\bm{\chi}}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|{\bm{\chi}}_{{\bm{u}}}\|_{0,\Omega}+\|{\bm{\chi}}_{\bm{\rho}}\|_{0,\Omega}+\|{\bm{\chi}}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}$
$\quad\leq C_{3}^{*}(\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\bm{\xi}_{{\bm{u}}}\|_{0,\Omega}+\|\bm{\xi}_{\bm{\rho}}\|_{0,\Omega}+\|\xi_{p}\|_{0,\Omega})$
$\qquad\displaystyle+\widetilde{C}^{*}_{5}\,(\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}+\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega})+\dfrac{1}{2}\|\chi_{p}\|_{0,\Omega}+\dfrac{1}{2}\|{\bm{\chi}}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}.$